OpenStudy (anonymous):

find the measures of two angles, one positive and one negative, that are coterminal with pie/7

OpenStudy (anonymous):

i got 15pie/7 ; -pie/7

OpenStudy (ghazi):

yes...i guess it's correct

OpenStudy (ghazi):

no the second one is incorrect

OpenStudy (ghazi):

it'll be 2 pi - pie/7

OpenStudy (anonymous):

thts not one of my choices though

terenzreignz (terenzreignz):

To get coterminal angles, just keep adding (or subtracting) 2pi until you get values that you want.

OpenStudy (anonymous):

ok then

OpenStudy (anonymous):

cause thats what i did

OpenStudy (anonymous):

do u guys want my choices ?

terenzreignz (terenzreignz):

Not with the second one, you didn't.... look again :)

terenzreignz (terenzreignz):

\[\frac{\pi}{7} - 2\pi\] ?

OpenStudy (ghazi):

to get positive terminal add 2 pie and to get negative terminal subtract 2 pie from pie/7

OpenStudy (ghazi):

so it'll be 15 pie/7 and -13 pie /7

OpenStudy (anonymous):

thats what i got in the beginning but the 13 didn't make sense

terenzreignz (terenzreignz):

Ok, why not? :) Want me to reassure you? lol

OpenStudy (ghazi):

why not?

terenzreignz (terenzreignz):

by the way, is -13pi/7 part of your choices?

OpenStudy (anonymous):

yes please

OpenStudy (anonymous):

it is thanks guys

OpenStudy (ghazi):

then it's 15 pie/7 and -13 pie/7

terenzreignz (terenzreignz):

No problem :)